OpenStudy (anonymous):

subtract: (3/(x-1))-((1-2x)/x)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

is the common denominator x(x-1) ?

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

ok I need help with the numerator. I know the first part is 3x-...??

OpenStudy (anonymous):

ok I get a little bunched up with the negative signs. I know the numerator is 3x-(1-2x)(x-1)

OpenStudy (anonymous):

could you help me multiply that? i get confused with the negatives.

OpenStudy (skullpatrol):

The best way to work with the negatives is to use a - b = a + (-b)

OpenStudy (skullpatrol):

$$\Huge 3x-(1+ (-2x))(x + (-1))$$

OpenStudy (skullpatrol):

Use FOIL

OpenStudy (anonymous):

do i multiply (1+(−2x))(x+(−1)) first? then distribute the negative?

OpenStudy (skullpatrol):

Yes

OpenStudy (anonymous):

ok let me do that. stay with me so you can check if my answer is right.

OpenStudy (anonymous):

3x−(1+(−2x))(x+(−1)) = 2x+1 ?

OpenStudy (skullpatrol):